OpenStudy (anonymous):

Given: Triangle ABC is similar to Triangle XYC. Enter answer as a simplified, improper fraction. If BY = 4, YC = 7, XC = 10, then AC = ? (angle drawn below)

OpenStudy (anonymous):



OpenStudy (anonymous):

@Desireegb shouldn't you have to separate the triangle to make it easier?

OpenStudy (cwrw238):

4/7 = AX / 10 AX = 40/7 AC = 10 + 40 / 7 = 110/7

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

your welcome

OpenStudy (cwrw238):

any more info needed?